jimmy sold 40 shirts for $150. if red shirts cost $3 each and blue shirt cost $4 each and jimmy only sold blue and red shirt,then how many of each shirt did he sell?

Answer by ikleyn(52782)

.

You can solve the problem using the system of 2 equations in 2 unknowns

From the condition, you have this system of 2 equation in 2 unknowns


    R +  B = 40              (1)    (counting red (R)  and blue (B) shirts)
   3R + 4B = 150  dollars    (2)    (counting dollars)


The most convenient way is to use Elimination method.
For it, multiply eq(1) by 3 (both sides) and then subtract the result from eq(2). You will get


    4B - 3B = 150 - 3*40,

     B = 150 - 120 = 30.


Answer.  30 blue shirts and the rest (40-30) = 10 red shirts.

Solved.   //   On the way you learned on how the Elimination method works.


Or you can solve it using one single equation.

Let B be the number of blue shirts.

Then the number of red shirts is  (40-B).


The money equation for the total is

    3*(40-B) + 4B = 150,

    120 - 3B + 4B = 150,

    B = 150 - 120 = 30,

and you get the same answer as in previous solution.